OpenStudy (anonymous):

Find the length of the longer leg of a 30°-60°-90° triangle with a hypotenuse of 6 in.

OpenStudy (anonymous):

Ratio of the sides of a right triangle are: 1 (short leg) 2 (hypotenuse) √3 (longer leg) Using ratios: hypotenuse / longer leg = hypotenuse / longer leg 2 / √3 = 6 / x Cross multiply: 2x = 6√3 x = 3√3 Using a calculator, that is about 5.2 inches.

OpenStudy (anonymous):

3in. 2\[\sqrt{3}\] 3\[\sqrt{3}\] 6\[\sqrt{3}\]

jhonyy9 (jhonyy9):

so do you know what sign sinus of an angle inside a right triangle ??

OpenStudy (anonymous):

wut?

jhonyy9 (jhonyy9):

so like sinA = opposite side divide hypothenus right ?

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

and we knoe again that to the smaler angle willé be oposite the shorter side yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

know wann to be there sorry

OpenStudy (anonymous):

its okay(:

jhonyy9 (jhonyy9):

so after these you van solve it sure right yes ?

OpenStudy (anonymous):

yes thank u

jhonyy9 (jhonyy9):

than good luck bye

jhonyy9 (jhonyy9):

than good luck bye